In an MRI, one detects the absorption and emission
Answer #1

E=hv

h is plank's constant and v is frequency

E=6.626 x 10-34*23*10^6

E=152.398*10^-28 J

a)E=gqhB/2m

2Em/ghq=B

2*152.398*10^-28*1.67*10^-27/2*6.626 x 10-34*1.6*10^-19

B=508.95*10^-55/21.2*10-53

B=24*10^-2 T

b) if g=2 then

E=gqhB/2m

E=2*1.6*10^-19*6.626 x 10-34*24*10^-2/2*(9.1*10^-31)

E=508.9*10^-55/18.2*10^-31

E=27.97*10^-24 J

E=hv

27.97*10^-24/6.626 x 10-34=v

4.23*10^10 Hz=v